OpenStudy (celinegirl):

CHECK MY ANSWER Evaluate cube root of 5 multiplied by square root of 5 over cube root of 5 to the power of 5. 5 to the power of negative 1 over 6 5 to the power of 3 over 2 5 to the power of 5 over 2 5 to the power of negative 5 over 6

OpenStudy (celinegirl):

Is the last one correct?

OpenStudy (anonymous):

\[\huge \frac{\sqrt[3]{5}\sqrt{5}}{\sqrt[3]{5^3}}\]?

OpenStudy (anonymous):

or maybe \[\huge \frac{\sqrt[3]{5}\sqrt{5}}{\sqrt[3]{5^5}}\]

OpenStudy (celinegirl):

Yes thats it

OpenStudy (anonymous):

same as \[\huge \frac{5^\frac{5}{6}}{5^{\frac{5}{3}}}\]

OpenStudy (anonymous):

no that is wrong

OpenStudy (anonymous):

\[\huge5^{-\frac{5}{6}}\]

OpenStudy (anonymous):

oh yea, last one

OpenStudy (celinegirl):

Oh okay! ha! I was getting a little confused. Thank you!

OpenStudy (celinegirl):

@satellite73 can you check one more? If you know graphs. i just wanna make sure i get a 100